How to solve this equation ?
2(5x-3)^(1/3) -4 = 0

Answer by lwsshak3(11628)

How to solve this equation ?
2(5x-3)^(1/3) -4 = 0
2(5x-3)^(1/3) =4
(5x-3)^(1/3) =4/2=2
cube both sides:
5x-3=2^3=8
5x-3=8
5x=11
x=11/5